Assume that 1-butene is being oxidized during a photochemical smog episode. Formulate the sequence of reactions that leads to the formation of two aldehyde species and using all peroxide radicals to oxidize NO· to NO2·.
Answer #1

formation of peroxide radical

H2O2 + O2 = 2HO2.

HO2. + NO = HO. + NO2

For the case of 1-Butene i.e. Alkene, HO . reacts through addition to either carbon atom of the double bond.

CH3CH2HC=CH2 + HO• → CH3CH2(HCOH)CH2

CH3CH2(HCOH)CH2 • + O2 → CH3CH2(HCOH)CH2O2 • , This is used to produce NO2

CH3(HCOH)CH2O2 • + NO → CH3(HCOH)CH2O• + NO2 .The net effect is to convert NO to NO2 and this produces more O3

CH3CH2CH=CH2 + O3 → CH3CH2CHOOOCH2 . This will fragment to produce two aldehyde species.

CH3CH2CHOOOCH2 → CH3CH2CHO + [• CH2OO• ] ≠

CH3CH2CHOOOCH2 → CH3CH2O + [CH3 • CHOO• ] ≠